# Quantum chaos and the arrow of time

### v10

@@ 4 Higher order terms of the change in energy
$\frac{i}{2}\sum_{l,m}E_{l}\left[\tilde{\lambda}(\omega_{nm})\tilde{\lambda}(\omega_{ml})\tilde{\lambda}(\omega_{ln})\mathcal{O}_{nm}\mathcal{O}_{ml}\mathcal{O}_{ln}-\tilde{\lambda}(\omega_{nl})\tilde{\lambda}(\omega_{lm})\tilde{\lambda}(\omega_{mn})\mathcal{O}_{nl}\mathcal{O}_{lm}\mathcal{O}_{mn}\right]$ (24)
@@
$\frac{1}{4}\,\sum_{k,m,l}(E_{n}+E_{l}-E_{k}-E_{m})[\tilde{\lambda}(\omega_{nk})\tilde{\lambda}(\omega_{kl})\tilde{\lambda}(\omega_{lm})\tilde{\lambda}(\omega_{mn})\mathcal{O}_{nk}\mathcal{O}_{kl}\mathcal{O}_{lm}\mathcal{O}_{mn}]$ (25)
In general, we find that this sum is negative (positive, zero) for initial states with positive (negative, infinite) temperature. $R_{mn}$ are not completely random [17]. If they were completely random, these higher order terms would be suppressed by exponentials of the entropy and a full analytic proof of the arrow of time at all orders of $\lambda(t)$ would be possible.
@@ 6 Conclusions
In a finite system, for which $f$ is not flat, the monotonic behaviour of $f(\bar{E})$ enhances or reinforces the arrow of time as we have seen in (23). For example, from the $\omega^{1}$ term of that equation, the system at positive temperature appears to absorb more energy per system size when the system size is small compared to a larger system, for same strength of the perturbation. It would be interesting to see if this result can be considered as a quantum advantage. We wish to study further in this direction in future.
It would also be interesting to check for ETH-monotonicity in long-range interacting systems for which equivalence of ensembles breaks down [31].
@@ Acknowledgements
The author would like to thank Gautam Mandal for helpful discussions during one of which the main idea for this paper arose. We also thank the referee 3 for an extremely thorough referee report with valuable insights.
@@
The author is fully supported by the Department of Science and Technology (Government of India) under the INSPIRE Faculty fellowship scheme
(Ref. No. DST/INSPIRE/04/2020/002105).
@@ Appendix A Change in energy starting from an energy eigenstate
The change in energy (7) is written in Schrodinger picture. In interaction picture, it can be written as
$\Delta E_{n}=\langle n|\bar{U}_{I}(t_{f})H_{0}U_{I}(t_{f})|n\rangle-\langle n|H_{0}|n\rangle$ (26)
@@
$\hskip 71.13188pt\mathcal{O}_{I}(t)=e^{i(t-t_{i})H_{0}}\mathcal{O}(t_{i})e^{-i(t-t_{i})H_{0}}$
$\bar{U}$ expansion is anti-time ordered and $U$ expansion is time-ordered. $\mathcal{O}_{I}(t)$ is the perturbing operator in interaction picture. With these expressions, the non-zero leading order term in (26) is
$\Delta E_{n}$ $=$ $\int_{t_{i}}^{t_{f}}dt_{1}dt^{\prime}_{1}\,\lambda(t_{1})\lambda(t^{\prime}_{1})\,\langle n|\mathcal{O}_{I}(t_{1})H_{0}\mathcal{O}_{I}(t^{\prime}_{1})|n\rangle$ (27)
$-\int_{t_{i}}^{t_{f}}dt_{1}\int_{t_{i}}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\,\langle n|\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})H_{0}+H_{0}\mathcal{O}_{I}(t_{1})\mathcal{O}_{I}(t_{2})|n\rangle$
@@
$\tilde{\lambda}(\omega)=\int_{-\infty}^{\infty}dt\,\lambda(t)e^{it\omega},\qquad\lambda(t)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\tilde{\lambda}(\omega)e^{-it\omega}$ (28)
@@
$\left.-E_{n}\int_{-\infty}^{\infty}dt_{1}\int_{-\infty}^{t_{1}}dt_{2}\int_{-\infty}^{\infty}\frac{d\omega_{1}d\omega_{2}}{(2\pi)^{2}}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega_{2})\left\{e^{it_{2}(E_{n}-E_{m}-\omega_{2})}e^{it_{1}(E_{m}-E_{n}-\omega_{1})}+(t_{1}\leftrightarrow t_{2})\right\}\right]$
@@
$-E_{n}\int_{-\infty}^{\infty}\frac{d\omega_{1}d\omega_{2}}{2\pi}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega_{2})\left\{\frac{\delta(\omega_{1}+\omega_{2})}{i(E_{n}-E_{m}-\omega_{2}-i\epsilon)}+\frac{\delta(\omega_{1}+\omega_{2})}{i(E_{m}-E_{n}-\omega_{2}-i\epsilon)}\right\}$ (29)
$=$ $-E_{n}|\tilde{\lambda}(E_{m}-E_{n})|^{2}$
Each term inside the parenthesis in the first line contributes half of the final result. The $\omega_{1}$ integrals are trivial. We remind here again that $\tilde{\lambda}(\omega)=\tilde{\lambda}(-\omega)^{*}$ because $\lambda(t)$ is a real function. The $\omega_{2}$ integrals are evaluated using Sokhotski–Plemelj theorem over the real line. The two Cauchy principal value integrals cancelled out because $|\tilde{\lambda}(\omega)|^{2}$ is an even function. Substituting these results, we get (8).
@@ Appendix B Change in energy starting from a thermal state
$\Delta E_{\beta}=\frac{1}{Z}\text{Tr}\left(U_{I}(t_{f})\rho_{0}\bar{U}_{I}(t_{f})\,H_{0}\right)-\frac{1}{Z}\text{Tr}\left(\rho_{0}H_{0}\right),\qquad Z=\text{Tr}\rho_{0}$ (30)
@@
$-\frac{1}{Z}\int_{t_{i}}^{t_{f}}dt_{1}\int_{t_{i}}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\;\sum_{n}\;\langle n|e^{-\beta H_{0}}\left(\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})H_{0}+H_{0}\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})\right)|n\rangle$
@@
$\Delta E_{\beta}$ $=$ $\frac{1}{Z}\int_{t_{i}}^{t_{f}}dt_{1}dt^{\prime}_{1}\,\lambda(t_{1})\lambda(t^{\prime}_{1})\left[\frac{1}{2}\,\sum_{n,m}E_{n}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}e^{it_{1}(E_{n}-E_{m})}e^{it^{\prime}_{1}(E_{m}-E_{n})}\right.$ (31)
@@
$\hskip 142.26378pt+\left.\sum_{n,m}E_{m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}e^{it_{1}(E_{m}-E_{n})}e^{it_{2}(E_{n}-E_{m})}\right]$
@@
$\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega_{1}d\omega^{\prime}_{1}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega^{\prime}_{1})\sum_{n,m}|\mathcal{O}_{nm}|^{2}\left[E_{n}e^{-\beta E_{m}}\delta(\omega_{1}-(E_{n}-E_{m}))\delta(\omega^{\prime}_{1}-(E_{m}-E_{n}))\right.$ (32)
@@
$=$ $\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega\,|\tilde{\lambda}(\omega)|^{2}\sum_{n,m}\left(E_{n}e^{-\beta E_{m}}+E_{m}e^{-\beta E_{n}}\right)|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))$
The integrals in the third and fourth lines in (31) are evaluated as in section A. The Cauchy principal values again cancelled out because $n$ and $m$ run over the entire state space, and $|\tilde{\lambda}(\omega)|^{2}$ is an even function. So, the third and fourth lines in (31) become
$-\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega\,|\tilde{\lambda}(\omega)|^{2}\sum_{n,m}\left(E_{n}e^{-\beta E_{n}}+E_{m}e^{-\beta E_{m}}\right)|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))$ (33)
@@
$=$ $\frac{1}{2}\int_{-\infty}^{\infty}d\omega\,\omega|\tilde{\lambda}(\omega)|^{2}\left[\frac{\left(e^{\beta\omega}-1\right)}{Z}\sum_{n,m}e^{\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))\right]$
This quantity is always a positive (negative or zero) at positive (negative or infinite) temperature $1/\beta$. The quantity inside the square brackets as a function of $\omega$ is the spectral function $A(\omega)$ [32, 33, 34] in a thermal state. [^6] The quantity inside the square brackets in (B) without the $\left(e^{\beta\omega}-1\right)$ factor is called spectral intensity in [32].
$A(\omega)$ $=$ $\frac{1}{Z}\left(e^{\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))$ (34)
$=$ $-\frac{1}{Z}\left(e^{-\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{n}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))$
@@ B.1 Derivation using linear response theory
To make the connection with linear response theory (LRT) clear, we rederive the change in energy using Kubo formula. We consider the time dependent Hamiltonian (6) where the real function $\lambda(t)$ has support for a finite range $-d<t<d$. The initial density matrix $\rho=e^{-\beta H_{0}}$ satisfies $[\rho,H_{0}]=0$. The rate of change of total energy [35] is given by
[…]
@@
$\langle\mathcal{O}(t)\rangle=\langle\mathcal{O}(0)\rangle+\int_{t_{i}}^{\infty}dt^{\prime}\,\lambda(t^{\prime})G_{R}(t,t^{\prime})$ (36)
$G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle$ (37)
$G_{R}$ is the definition of retarded Green function (in general cases, also called response function) of the operator $\mathcal{O}$. This relation is also known as the Kubo formula for the response function [36, 35]. Considering $t_{i}<-d$ and $d<t_{f}$, the total energy gained upto this order is
$\Delta E_{\beta}=\int_{t_{i}}^{t_{f}}dt\,\frac{dE}{dt}=\langle\mathcal{O}(0)\rangle\int_{t_{i}}^{t_{f}}dt\,\dot{\lambda}(t)+\int_{t_{i}}^{t_{f}}dt\,\dot{\lambda}(t)\int_{t_{i}}^{\infty}dt^{\prime}\,G_{R}(t,t^{\prime})\lambda(t^{\prime})$ (38)
@@
$\Delta E_{\beta}=i\int_{-\infty}^{\infty}d\omega\omega|\lambda(\omega)|^{2}\;G_{R}\left(\omega\right)=\frac{1}{2}\int_{-\infty}^{\infty}d\omega\omega|\lambda(\omega)|^{2}\;A\left(\omega\right)$ (39)
The real part of $G_{R}(\omega)$ is even and the imaginary part of $G_{R}(\omega)$ is odd and $A(\omega)=-2\,\text{Im}\;\tilde{G}_{R}(\omega)$.

### v12

@@ 4 Higher order terms of the change in energy
$\frac{i}{2}\sum_{l,m}E_{l}\left[\tilde{\lambda}(\omega_{nm})\tilde{\lambda}(\omega_{ml})\tilde{\lambda}(\omega_{ln})\mathcal{O}_{nm}\mathcal{O}_{ml}\mathcal{O}_{ln}-\tilde{\lambda}(\omega_{nl})\tilde{\lambda}(\omega_{lm})\tilde{\lambda}(\omega_{mn})\mathcal{O}_{nl}\mathcal{O}_{lm}\mathcal{O}_{mn}\right]\$(4.1)
@@
$\frac{1}{4}\,\sum_{k,m,l}(E_{n}+E_{l}-E_{k}-E_{m})[\tilde{\lambda}(\omega_{nk})\tilde{\lambda}(\omega_{kl})\tilde{\lambda}(\omega_{lm})\tilde{\lambda}(\omega_{mn})\mathcal{O}_{nk}\mathcal{O}_{kl}\mathcal{O}_{lm}\mathcal{O}_{mn}]\$(4.2)
In general, we find that this sum is negative (positive, zero) for initial states with positive (negative, infinite) temperature. $R_{mn}$ are not completely random Foini:2018sdb . If they were completely random, these higher order terms would be suppressed by exponentials of the entropy and a full analytic proof of the arrow of time at all orders of $\lambda(t)$ would be possible.
@@ 6 Conclusions
In a finite system, for which $f(\bar{E})$ is not flat, the monotonic behaviour of $f(\bar{E})$ enhances or reinforces the arrow of time as we have seen in (3.19). For example, from the $\omega^{1}$ term of that equation, the system at positive temperature appears to absorb more energy per system size when the system size is small compared to a larger system, for same strength of the perturbation. It would be interesting to see if this result can be considered as a quantum advantage. We wish to study further in this direction in future.
It would also be interesting to check for ETH-monotonicity in long-range interacting systems for which equivalence of ensembles breaks down Barr__2001 .
@@ Acknowledgements
The author would like to thank Gautam Mandal for helpful discussions during one of which the main idea for this paper arose.
@@
The author is fully supported by the Department of Science and Technology (Government of India) under the INSPIRE Faculty fellowship scheme (IFA-20-PH-262).
@@ Appendix A Change in energy starting from an energy eigenstate
The change in energy (3.2) is written in Schrodinger picture. In interaction picture, it can be written as
$\Delta E_{n}=\langle n|\bar{U}_{I}(t_{f})H_{0}U_{I}(t_{f})|n\rangle-\langle n|H_{0}|n\rangle\$(A.1)
@@
$\hskip 71.13188pt\mathcal{O}_{I}(t)=e^{i(t-t_{i})H_{0}}\mathcal{O}(t_{i})e^{-i(t-t_{i})H_{0}}\$
$\bar{U}$ expansion is anti-time ordered and $U$ expansion is time-ordered. $\mathcal{O}_{I}(t)$ is the perturbing operator in interaction picture. With these expressions, the non-zero leading order term in (A.1) is
$\Delta E_{n}$ $=$ $\int_{t_{i}}^{t_{f}}dt_{1}dt^{\prime}_{1}\,\lambda(t_{1})\lambda(t^{\prime}_{1})\,\langle n|\mathcal{O}_{I}(t_{1})H_{0}\mathcal{O}_{I}(t^{\prime}_{1})|n\rangle$ (A.2)
$-\int_{t_{i}}^{t_{f}}dt_{1}\int_{t_{i}}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\,\langle n|\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})H_{0}+H_{0}\mathcal{O}_{I}(t_{1})\mathcal{O}_{I}(t_{2})|n\rangle\$
@@
$\tilde{\lambda}(\omega)=\int_{-\infty}^{\infty}dt\,\lambda(t)e^{it\omega},\qquad\lambda(t)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\tilde{\lambda}(\omega)e^{-it\omega}\$(A.3)
@@
$\left.-E_{n}\int_{-\infty}^{\infty}dt_{1}\int_{-\infty}^{t_{1}}dt_{2}\int_{-\infty}^{\infty}\frac{d\omega_{1}d\omega_{2}}{(2\pi)^{2}}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega_{2})\left\{e^{it_{2}(E_{n}-E_{m}-\omega_{2})}e^{it_{1}(E_{m}-E_{n}-\omega_{1})}+(t_{1}\leftrightarrow t_{2})\right\}\right]\$
@@
$-E_{n}\int_{-\infty}^{\infty}\frac{d\omega_{1}d\omega_{2}}{2\pi}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega_{2})\left\{\frac{\delta(\omega_{1}+\omega_{2})}{i(E_{n}-E_{m}-\omega_{2}-i\epsilon)}+\frac{\delta(\omega_{1}+\omega_{2})}{i(E_{m}-E_{n}-\omega_{2}-i\epsilon)}\right\}$ (A.4)
$=$ $-E_{n}|\tilde{\lambda}(E_{m}-E_{n})|^{2}\$
Each term inside the parenthesis in the first line contributes half of the final result. The $\omega_{1}$ integrals are trivial. We remind here again that $\tilde{\lambda}(\omega)=\tilde{\lambda}(-\omega)^{*}$ because $\lambda(t)$ is a real function. The $\omega_{2}$ integrals are evaluated using Sokhotski–Plemelj theorem over the real line. The two Cauchy principal value integrals cancelled out because $|\tilde{\lambda}(\omega)|^{2}$ is an even function. Substituting these results, we get (3.3).
@@ Appendix B Change in energy starting from a thermal state
$\Delta E_{\beta}=\frac{1}{Z}\text{Tr}\left(U_{I}(t_{f})\rho_{0}\bar{U}_{I}(t_{f})\,H_{0}\right)-\frac{1}{Z}\text{Tr}\left(\rho_{0}H_{0}\right),\qquad Z=\text{Tr}\rho_{0}\$(B.1)
@@
$-\frac{1}{Z}\int_{t_{i}}^{t_{f}}dt_{1}\int_{t_{i}}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\;\sum_{n}\;\langle n|e^{-\beta H_{0}}\left(\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})H_{0}+H_{0}\mathcal{O}_{I}(t_{2})\mathcal{O}_{I}(t_{1})\right)|n\rangle\$
@@
$\Delta E_{\beta}$ $=$ $\frac{1}{Z}\int_{t_{i}}^{t_{f}}dt_{1}dt^{\prime}_{1}\,\lambda(t_{1})\lambda(t^{\prime}_{1})\left[\frac{1}{2}\,\sum_{n,m}E_{n}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}e^{it_{1}(E_{n}-E_{m})}e^{it^{\prime}_{1}(E_{m}-E_{n})}\right$. (B.2)
@@
$\hskip 142.26378pt+\left.\sum_{n,m}E_{m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}e^{it_{1}(E_{m}-E_{n})}e^{it_{2}(E_{n}-E_{m})}\right]\$
@@
$\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega_{1}d\omega^{\prime}_{1}\,\tilde{\lambda}(\omega_{1})\tilde{\lambda}(\omega^{\prime}_{1})\sum_{n,m}|\mathcal{O}_{nm}|^{2}\left[E_{n}e^{-\beta E_{m}}\delta(\omega_{1}-(E_{n}-E_{m}))\delta(\omega^{\prime}_{1}-(E_{m}-E_{n}))\right$. (B.3)
@@
$=$ $\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega\,|\tilde{\lambda}(\omega)|^{2}\sum_{n,m}\left(E_{n}e^{-\beta E_{m}}+E_{m}e^{-\beta E_{n}}\right)|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))\$
The integrals in the third and fourth lines in (B.2) are evaluated as in section A. The Cauchy principal values again cancelled out because $n$ and $m$ run over the entire state space, and $|\tilde{\lambda}(\omega)|^{2}$ is an even function. So, the third and fourth lines in (B.2) become
$-\frac{1}{2Z}\int_{-\infty}^{\infty}d\omega\,|\tilde{\lambda}(\omega)|^{2}\sum_{n,m}\left(E_{n}e^{-\beta E_{n}}+E_{m}e^{-\beta E_{m}}\right)|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))\$(B.4)
@@
$=$ $\frac{1}{2}\int_{-\infty}^{\infty}d\omega\,\omega|\tilde{\lambda}(\omega)|^{2}\left[\frac{\left(e^{\beta\omega}-1\right)}{Z}\sum_{n,m}e^{\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))\right]\$
This quantity is always a positive (negative or zero) at positive (negative or infinite) temperature $1/\beta$. The quantity inside the square brackets as a function of $\omega$ is the spectral function $A(\omega)$ Zubarev_1960 ; sglrbook2013 ; coleman15 in a thermal state. [^5] The quantity inside the square brackets in (B) without the $\left(e^{\beta\omega}-1\right)$ factor is called spectral intensity in Zubarev_1960 .
$A(\omega)$ $=$ $\frac{1}{Z}\left(e^{\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))$ (B.5)
$=$ $-\frac{1}{Z}\left(e^{-\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{n}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n}))\$
@@ B.1 Derivation using linear response theory
To make the connection with linear response theory (LRT) clear, we rederive the change in energy using Kubo formula. We consider the time dependent Hamiltonian (3.1) where the real function $\lambda(t)$ has support for a finite range d<t<d-d<t<d. The initial density matrix $\rho=e^{-\beta H_{0}}$ satisfies $\rho,H_{0}]=0$. The rate of change of total energy jjmrbook1991 is given by
$\frac{dE}{dt}=\frac{d}{dt}\langle\mathcal{H}\rangle=\left\langle\frac{d\mathcal{H}}{dt}\right\rangle=\dot{\lambda}(t)\langle\mathcal{O}(t)\rangle\$(B.6)
@@
$\langle\mathcal{O}(t)\rangle=\langle\mathcal{O}(0)\rangle+\int_{t_{i}}^{\infty}dt^{\prime}\,\lambda(t^{\prime})G_{R}(t,t^{\prime})$ (B.7)
$G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle\$(B.8)
$G_{R}$ is the definition of retarded Green function (in general cases, also called response function) of the operator $\mathcal{O}$. This relation is also known as the Kubo formula for the response function Kubo:1957mj ; jjmrbook1991 . Considering $t_{i}<-d$ and $d<t_{f}$, the total energy gained upto this order is
$\Delta E_{\beta}=\int_{t_{i}}^{t_{f}}dt\,\frac{dE}{dt}=\langle\mathcal{O}(0)\rangle\int_{t_{i}}^{t_{f}}dt\,\dot{\lambda}(t)+\int_{t_{i}}^{t_{f}}dt\,\dot{\lambda}(t)\int_{t_{i}}^{\infty}dt^{\prime}\,G_{R}(t,t^{\prime})\lambda(t^{\prime})\$(B.9)
@@
$\Delta E_{\beta}=i\int_{-\infty}^{\infty}d\omega\omega|\lambda(\omega)|^{2}\;\tilde{G}_{R}\left(\omega\right)=\frac{1}{2}\int_{-\infty}^{\infty}d\omega\omega|\tilde{\lambda}(\omega)|^{2}\;A\left(\omega\right)\$(B.10)
The real part of $\tilde{G}_{R}(\omega)$ is even and the imaginary part of $\tilde{G}_{R}(\omega)$ is odd and $A(\omega)=-2\,\text{Im}\;\tilde{G}_{R}(\omega)$.